OpenStudy (anonymous):

Write the equation in slope intercept form of the line parallel to -3x+2y=9 and goes through the point (-2,1)

OpenStudy (anonymous):

so it is parallel to -3x+2y=9, hence they have the same slope. we can take the slope from y=3/2x+9/2, where m=3/2. then just plug into (y-y0)=m(x-x0) to find the equation of the line, where x0=-2 and y0=1 (in case you aren't used to the notation) and you will get: y=3/2x-2

OpenStudy (anonymous):

m = 3/2 y-1 = 3/2(x+2)

OpenStudy (anonymous):

http://answers.yahoo.com/question/index?qid=20101231194014AATaDoA this explains it.

OpenStudy (anonymous):

y = 3/2x +4

OpenStudy (anonymous):

^ that is the right answer, i made a sign mistake in my asnwer.

OpenStudy (anonymous):

@Directrix